# Lifelong Sequential Knowledge Editing without Model Degradation

### v1

@@ 4 Overfitting during Knowledge Editing
As discussed in section 2, locate-then-edit methods can be seen as a two-step fine tuning process. The first step uses gradient descent to create intermediate target representations. The gradient descent step minimizes average cross-entropy loss for predicting the target fact over the query phrase augmented by ‘$N$’ random prefixes. The random prefixes are supposed to represent different contexts in which the edited fact can be recalled, thus aiming to create a more general query representation. The average cross-entropy loss over all contexts is optimized as shown:
$L(\theta)=\frac{1}{N}\sum^{j=N}_{j=1}-log\mathcal{P}_{\theta}[o^{*}|x_{j}+p]$ (2)
Here $p$ represents the query phrase input to the model, $o^{*}$ represents the target fact to be edited into the model and $x_{j}$ represent random prefixes added to the query phrase.
(a) MEMIT
(b) AlphaEdit
(c) EMMET
Figure 3: Average downstream performance measured over 6 tasks (sec 2) for MEMIT, AlphaEdit and EMMET. We see that MPES is able to delay loss of downstream performance for both Llama3-8B and Llama2-7B with additional gains in efficiencey.
Table 2: Sequential knowledge editing performance after 10,000 edits for different algorithms. We see that using MPES improves editing metrics across all algorithms and models.
@@
Method Model Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy
MEMIT Llama2-7B 81.04 64.67 60.95 67.859 442.59
Llama3-8B 49.68 49.29 51.31 50.078 373.48
MEMIT Llama2-7B 88.43 70.83 65.86 73.873 542.1
+ MPES Llama3-8B 65.78 57.58 50.25 57.176 560.78
AlphaEdit Llama2-7B 61.1 55.86 53.75 56.74 540.92
Llama3-8B 72.67 63.44 52.9 61.948 465.81
AlphaEdit Llama2-7B 84.15 74.94 62.87 72.933 583.4
+ MPES Llama3-8B 88.43 82.08 56.5 72.832 565.36
EMMET Llama2-7B 94.98 84.05 55.76 74.331 569.24
Llama3-8B 85.03 75.06 49.08 65.995 567.34
EMMET Llama2-7B 96.8 88.4 58.41 77.393 584.28
+ MPES Llama3-8B 92.96 86.12 55.74 74.424 564.82
@@
We observe that when the gradient descent process stops, the intermediate target representations are over-optimized on the edited facts and make the probability of the target fact unusually high. This can be seen in Table 1, where the edited facts are predicted with a probability much closer to 1 compared to the probability with which an unedited model recalls a fact. This phenomenon was also observed by Zhang et al. (2024). More details on the experimental settings can be found in Appendix B.
As seen in Table 1, the probability with which the unedited GPT2-XL, Llama2-7B, and Llama3-8B models recall a fact are 39%, 52% and 49% respectively. However, facts that get edited into the model with algorithms like Alphaedit or EMMET are predicted with average probabilities of 98-99%. The vocabulary size of these models is in the tens of thousands and placing such a high probability mass over every edited fact is a result of clear overfitting. We hypothesize that this overfitting of edited facts is one of the reasons why edited models lose their ability to perform downstream tasks when continuously edited using knowledge editing methods.
@@ 4.1 MPES and Knowledge Editing
To overcome this overfitting over a small subset of edited facts, we propose a variant of early stopping called “most-probable early stopping” (MPES). Conventionally, early stopping is used during training while monitoring validation loss, where training is halted when the validation loss stops improving. In MPES, we stop the gradient descent process in knowledge editing when the target fact becomes the most probable token for all ‘$q$’ query phrases used for optimization. While the gradient descent loss continues to go down and the target probability continues to increase, the rank of the predicted word does not change further. With MPES, the rank of the predicted word performs the same function as validation loss in conventional early stopping.
@@
Apart from preventing the model from becoming overly optimized towards a specific target fact, using MPES for halting gradient descent also has two other advantages. Firstly, it simplifies monitoring of the gradient descent process and provides a principled approach to stopping gradient descent which is directly tied to the knowledge editing objective, that is, accurately recalling edited facts in a variety of scenarios without overfitting. Secondly, by optimally stopping the gradient descent process, we also improve the efficiency of locate-then-edit algorithms. MPES leads to a 39-76% reduction in editing time per edit depending on the editing algorithm and model used (Table 1).
The average downstream performance of AlphaEdit, MEMIT, and EMMET for Llama2-7B and Llama3-8B over 10,000 sequential edits is shown in Figure 3111For task-wise downstream performance, please refer to Appendix H. Comparing baseline methods with MPES, we see that MPES shows significant improvements in downstream performance for both AlphaEdit and MEMIT. For EMMET, MPES performs on par with the baseline but is much faster, as shown in Table 1. This shows that by halting gradient descent at a critical point where the edited fact is recognized but not overly dominant, MPES helps maintain the model’s general abilities for much longer during sequential editing.
The editing performance using MPES can be seen in Table 2. We see that MPES not only maintains the model’s general abilities for longer but also improves editing performance for all algorithms. This is because as the model over-optimizes for edited facts, it not only loses its general abilities but also the ability to recall facts edited in previous iterations, resulting in lower editing scores. This also verifies our hypothesis that overfitting over edited facts is one of the reasons behind the loss of model’s general ability.
To summarize, with MPES we present a principled way of stopping the gradient descent step during knowledge editing which results in improved editing performance, delays the loss of downstream performance, and makes current knowledge editing methods much faster.
(a) Alphaedit (Llama3-8B)
(b) MEMIT (Llama3-8B)
Figure 4: Comparison between norm of edited MLP matrices and norm of unedited matrices after 5,000 and 10,000 sequential edits.
@@ 5 Norm Growth during Sequential Knowledge Editing
Past work has shown that sequential knowledge editing can lead to an increase in the norm of the edited matrix (Gupta et al., 2024b). This disproportionate growth of norm can be seen in Figure 1 and 4 when using AlphaEdit and MEMIT on Llama3-8B. Both AlphaEdit and MEMIT modify the second MLP matrix of layers 4 to 8. To put the norm growth in perspective with the rest of the model, we plot the norms of the edited matrices after 5k and 10k edits along with the norm of other MLP matrices of the model. We clearly see the disproportionate growth in the norm for the edited layers in Figure 4, where the norm grows to more than 10 times its original value when using MEMIT, and twice the original norm when using AlphaEdit, while the norm of the matrices before and after the edited layers is not changed. The growth of the norm of the edited matrix is a continuous process during knowledge editing. As shown in Figure 1, for not even one edit does the norm of the edited matrix remain constant or decrease during editing. The norm always increases!
But is this growing norm harmful to the model? We answer this question by analyzing the residual stream and reveal a very important property about the inner workings of knowledge editing methods.
5.1 The Secret Mechanics of Knowledge Editing
Due to residual connections in an LLM, the final output of a model can be written as a summation of the outputs of the individual sub-modules. Let $h^{l}$ represent the intermediate hidden state vectors between each decoder layer. Then, the computations within a layer of a general decoder-only LLMs proceed as follows:
[…]
$a^{l}$ $=\texttt{Att}(f^{l})$ (4)
[…]
$m^{l}$ $=W^{l}_{proj}\sigma(W^{l}_{fc}g^{l}+b^{l}_{fc})+b_{proj}$ (6)
$h^{l}$ $=h^{l-1}+a^{l}+m^{l}$ (7)
The intermediate hidden vector between each layer, $h^{l}$, is also sometimes referred to as the residual stream. LN1 represents the first LayerNorm (or equivalently RMSNorm for Llama models) that acts just before the attention module and LN2 represents the second LayerNorm just before the MLP module. Att represents the self-attention module in an LLM whereas the action of a traditional MLP module is written in terms of the individual weight matrices ($W_{fc},W_{proj}$). As the vectors computed in the attention and MLP modules get added back to the residual stream at each layer, the residual stream represents a summation of an increasing number of vectors as we go deeper into the model. A non-recursive formula for the output of the transformer just before unembedding is shown below:
@@ 5.1 The Secret Mechanics of Knowledge Editing
Here, $L$ represents the total number of layers in a model and $h^{L}$ represents the residual stream after the final layer. Note that $h^{L}$ is the activation just before the application of the final layernorm and unembedding matrices. Thus, the output vector at the final layer is a summation of the outputs of individual self-attention and MLP sub-modules.
@@
Now, if the norm of the $W_{proj}$ matrix grows as disproportionately as shown in Figures 1 and 4, the norm of the vectors that are produced from those edited MLP sub-modules will also grow. As the norm of a few vectors in the summation in equation 8 grows, these vectors will begin to dominate the sum. Proof for this is shown in Appendix C.1, where we show that if the norm of a vector in a summation grows, the overall sum effectively tends towards that vector.
(a) Average Norm Proportion For Unedited Llama3-8B
(b) Average Norm Proportion for Llama3-8B using Alphaedit
(c) Average Norm Proportion for Llama3-8B using MEMIT
(d) Average Norm Proportion for Llama3-8B using ENCORE
Figure 5: The figure shows the proportion of contribution of activations vectors produced from each sub-module to the residual stream. The edited layers are shown in red. We see that the influence of the output of the edited layers grows dramatically after contiuous editing.
@@
We also show this effect empirically. The growing norm of activation vectors produced by edited layers before and after editing for Llama3-8B can be seen in Figure 5. Figure 5(a) shows the percentage norm for activation vectors produced from each sub-module of the unedited Llama3-8B model, in comparison to the sum of their norms [^2] This is averaged over 30,000 generated tokens (Appendix C).. After editing using AlphaEdit, which edits layers 4-8 in the Llama model for 10,000 sequential edits (edited layers are shown in red color on x-axis), we can see a drastic increase in the percentage norms for vectors produced by the edited layers (Figure 5(b)). For example, the norm of vectors produced by layer 8, which accounted for only 1% of the total norm, after editing using AlphaEdit accounts for 7% of the norm. This effect is even more drastic for MEMIT as seen in Figure 5(c), where the activation vectors produced by layer 8 account for almost 40% of the total norm, and the vectors produced by all edited layers account for 85% of the total.
@@
This gives us a crucial insight into the inner workings of locate-then-edit methods. Apart from optimizing for the preservation-memorization objective in equation 1, these methods also implicitly increase the norms of the edited matrices, and as a consequence, the norm of the activation vectors produced from those matrices also increases. This makes the representation after the final layer ($h^{L}$) be dominated by the output of edited layers, thus giving the edited layers a larger authority over the final representations. This allows edited layers to override the information produced from other parts of the model, consequently leading to the success of these knowledge editing methods.
This finding is central to explaining the inner workings of locate-then-edit methods, and it also explains their successes and failures. This kind of “importance hacking” makes the output of the edited layers more dominant in the final output and is a shortcut taken by these methods to achieve edit success. However, as the norm of the edited matrix continues to grow when we edit it sequentially, suddenly the final representations begin to largely be composed of the outputs of only the edited layers. This makes it impossible for the model to account for the information processed from other sub-modules which might be important to perform other downstream or unrelated tasks. We hypothesize that this is another cause of the loss of downstream performance.
Table 3: Editing performance of ENCORE when compared to baseline MEMIT and modifications using MPES and Norm Constraint.
Method Model Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy
MEMIT Llama2-7B 81.04 64.67 60.95 67.859 442.59
Llama3-8B 49.68 49.29 51.31 50.078 373.48
MEMIT Llama2-7B 88.43 70.83 65.86 73.873 542.1
+ MPES Llama3-8B 65.78 57.58 50.25 57.17 560.78
MEMIT Llama2-7B 90.94 81.31 59.73 74.931 539.58
+ Norm-Const Llama3-8B 85.72 77.08 58.45 71.86 367.46
ENCORE Llama2-7B 92.57 82.64 60.43 76.043 560.16
Llama3-8B 88.77 78.19 60.07 73.707 523.61
5.2 Introducing ENCORE
@@
In the above discussion, we show how growing norm of edited matrices is detrimental to the functioning of edited models and hypothesize that it is the cause of loss of downstream performance. To test our hypothesis, we propose to add an additional term to the preservation-memorization objective to control this norm growth. Thus, we augment the MEMIT objective with a norm-constraint:
@@
The original weight matrix is represented by $W_{0}$ and $\hat{W}$ represents the edited matrix. The above objective has a closed-form solution as shown below (proof, Appendix D):
@@
Editing performance for 10,000 sequential edits using norm-constraint can be seen in Table 3. Since we add the norm-constraint to the MEMIT objective  [^3] While adding norm-constraint works well with the MEMIT objective, using norm constraint with EMMET and AlphaEdit does not lead to such extensive improvements. We discuss this further in Appendix E., we compare the performance of MEMIT and MEMIT+MPES. We see a significant improvement in all editing metrics for both Llama2-7B and Llama3-8B. The downstream performance can be seen in Figure 6, where norm-constraint sustains downstream performance for much longer compared to MEMIT and MPES.
@@
Figure 6: Average downstream performance for during sequential editing with ENCORE compared to baseline of MEMIT and addition of MPES and Norm-Constraint (NC).
(a) Comparison with unedited layers.
(b) Norm growth as function of edits.
Figure 7: Growth in norm of edited matrices for ENCORE for Llama3-8B.
To achieve our goal of performing 10,000 sequential edits without loss of downstream performance, we combine MPES with the norm-constraint objective in equation 9 and present ENCORE - Early stopping and Norm-Constrained Robust knowledge Editing. With ENCORE, we bring back the classical knowledge editing objective of MEMIT in a stronger, faster, and more robust way. The editing performance for ENCORE can be seen in Table 3 and Figure 6. For Llama2-7B, we see that maximum improvement in downstream performance (Figure 6(a)) comes from the norm constraint and ENCORE performs equally well. But ENCORE is not only much faster, but also has a better overall editing performance (Table 3). For Llama3-8B (Figure 6(b)), we see that the combination of MPES and norm-constraint is crucial for sustained downstream performance.
ENCORE also controls the growing norm of edited matrices, as shown in Figure 7. As a consequence, the activation norms for ENCORE have a very similar distribution to the original model, which is shown in Figure 5(d). Thus, ENCORE is able to perform large-scale knowledge editing without taking the shortcut of importance hacking, which also prevents the loss of downstream performance. With ENCORE, we combine the insights presented in this paper about overfitting and implicit importance hacking in locate-then-edit methods and enable long-term sequential knowledge editing. ENCORE is able to consistently sustain downstream evaluation metrics of the original model while going through 10,000 sequential edits. ENCORE is also 61% faster than MEMIT and 64% faster than AlphaEdit for Llama3-8B [^4] ENCORE takes 1.01 seconds per edit for GPT2-XL, 3.33 seconds for Llama2-7B and 3.35 seconds for Llama3-8B..

### v2

@@ 5 Over-Optimization of Target Activations
The gradient-descent step for finding target activations as described in section 4 minimizes average cross-entropy loss for predicting the target fact for the query phrase augmented by ‘$N$’ random prefixes. The random prefixes are supposed to represent different contexts in which the edited fact can be recalled, thus aiming to create a more general query representation Meng et al. (2022a). Let $p$ represent the query-phrase input to the model, $o^{*}$ represent the target fact to be edited into the model and $x_{j}$ represent random prefixes added to the query phrase. Then, the gradient-descent step minimizes the following loss function:
$L(\theta)=\frac{1}{N}\sum^{j=N}_{j=1}-log\mathcal{P}_{\theta}[o^{*}|x_{j}+p],$ (2)
@@
Method Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy Llama2-7B Llama3-8B Llama2-7B Llama3-8B Llama2-7B Llama3-8B Llama2-7B Llama3-8B Llama2-7B Llama3-8B MEMIT 81.04 49.68 64.67 49.29 60.95 51.31 67.86 50.08 442.59 373.48 MEMIT + LTI 55.34 51.85 53.57 51.23 49.21 49.21 52.58 50.70 562.07 372.08 MEMIT + MPES 93.21 74.40 83.43 66.32 62.25 50.78 77.36 62.23 569.84 466.97 RECT 82.42 63.17 66.84 56.92 67.39 52.89 71.54 57.36 549.35 588.39 RECT + LTI 53.87 51.55 52.17 51.08 50.35 50.43 52.09 51.02 498.97 544.41 RECT + MPES 92.12 66.52 78.06 59.11 67.16 59.44 77.81 61.51 596.62 410.82 PRUNE 70.80 49.38 62.11 49.63 51.86 51.09 60.60 50.02 280.83 340.22 PRUNE + LTI 55.19 49.88 52.98 49.63 50.33 50.19 52.76 49.90 548.23 239.17 PRUNE + MPES 91.48 92.66 83.65 82.20 61.91 64.82 76.85 78.16 568.91 512.83 AlphaEdit 61.10 72.67 55.86 63.44 53.75 52.90 56.74 61.95 540.92 465.81 AlphaEdit + LTI 53.82 53.27 53.03 51.70 49.68 49.72 52.11 51.52 524.11 240.07 AlphaEdit + MPES 84.15 88.43 74.94 82.08 62.87 56.60 72.93 72.83 583.40 565.32
Table 2: Sequential knowledge-editing performance after 10,000 edits. Scores are shown for nine MEMIT-based and AlphaEdit algorithms under two models. MPES consistently improves editing metrics across all settings.
@@
We observe that when the gradient-descent step stops in these algorithms, the intermediate target representations become over-optimized, assigning unusually high probabilities to edited facts. More specifically, the average target probability at which the gradient-descent step stops for all locate-then-end algorithms lies between 95-100%. When this target representation is used to update the the edited matrix using equation 1, the resultant edits are predicted with abnormally high probabilities. This can be seen in Table 1, where the edited facts are predicted with a much larger probability compared to the original “unedited” models. More details on the experimental settings can be found in Appendix B. This phenomenon was also observed by Zhang et al. (2024b).
@@ 5.1 MPES and Knowledge Editing
As seen in Table 1, the probability with which the unedited Llama2-7B, and Llama3-8B naturally recall a fact are 52% and 49% respectively. However, facts that get edited into the model with algorithms like MEMIT are predicted with an average probability of 78-79%. To overcome this hyper-optimization over a small subset of edited facts, we propose a variant of early stopping called “most-probable early stopping" (MPES).
@@
Conventionally, early stopping is used during training while monitoring validation loss, where training is halted when the validation loss stops improving. In MPES, we stop the gradient-descent step in knowledge editing when the target fact becomes the most probable token for all ‘$N$’ query phrases used for optimization (equation 2). This contrasts with the current stopping criteria in these methods, where gradient-descent is halted after either 20 iterations or upon reaching a loss threshold. These criteria leave open the possibility of over-optimization (and hence overfitting, as observed empirically in Table 1), or underfitting. Apart from preventing the model from becoming overly optimized towards a specific target fact, using MPES for halting gradient-descent also has two other advantages. Firstly, it simplifies monitoring of the gradient-descent process and provides a principled approach to stopping gradient-descent which is directly tied to the knowledge editing objective, that is, accurately recalling edited facts in a variety of scenarios without overfitting. Secondly, by optimally stopping the gradient-descent process, we also improve the efficiency of locate-then-edit algorithms. MPES reduces the average edit time by 42-61% compared to the standard MEMIT.
(a) Llama2-7B
(b) Llama3-8B
Figure 3: Downstream evaluation when comparing MPES (our method) with LTI for regularizing over-optimization of target activations during knowledge editing.
We compare MPES with LTI Zhang et al. (2024b), which prevents over-optimization during the gradient-descent step using multiple additional loss functions that also increase the average time per edit (LTI takes approximately three times as long as MPES as shown in Table 1). We combine MPES with multiple knowledge editing methods including MEMIT Meng et al. (2022b), RECT Gu et al. (2024b), PRUNE Ma et al. (2025) and AlphaEdit Fang et al. (2024). We perform 10,000 sequential edits in batches of 100 as described in section 3. The editing and downstream evaluation metrics for sequential editing for both regularization methods can be seen in Table 2 and Figure 3 respectively. We see that MPES provides substantial gain across all key metrics.
Table 2 shows that all editing metrics are enhance significantly for all knowledge editing methods when combined with MPES. Infact, Table 2 shows that LTI harms the editing metrics for sequential editing when compared to even the respective baselines. The downstream performance metrics can be seen in Figure 3. We see that knowledge editing methods augmented with MPES (represented by solid lines with diamond dots) are able to maintain their downstream performance for much longer when compared to the corresponding baselines and LTI. However, despite these improvements, downstream degradation still occurs after a few thousand edits especially in larger models like Llama3-8B. This suggests that MPES alone may not fully address all failure modes, which we explore further in the next section. To summarize, with MPES we present a principled way of stopping the gradient-descent step during knowledge editing which results in improved editing performance, delays loss of downstream performance, and makes current knowledge editing methods much faster.
@@ 6 Norm Growth during Sequential Knowledge Editing
Previous studies show that sequential knowledge editing increases the norm of the edited matrix Gupta et al. (2024b); Yang et al. (2024); Ma et al. (2025). Figure 1 illustrates this issue, highlighting the extreme norm growth in edited layers of Llama3-8B using MEMIT. Here, the norms of the edited layers increase by more than 10 times, while the unedited layer remains unchanged (Figure 1(b)). Furthermore, norm growth persists continuously during editing, as shown in Figure 1, for not even one edit does the norm remain constant or decrease.
While the anomalous norm-growth was observed in prior work, they do not explain how it affects the general ability of the model. We answer this question by analyzing the residual stream of the model during large-scale knowledge editing.
6.1 Explaining Loss of Downstream Performance due to Norm-Growth
To understand the impact of this norm growth, we analyze how residual connections work in decoder-only LLMs. The intermediate hidden vector at layer $l$, represented by $h^{l}$, is also sometimes referred to as the residual stream. Each decoder-only layer contains one attention and FFN module that feeds directly into the residual stream through residual connections. The exact computations happening within transformer-based decoder-only LLMs can be found in Appendix E. Let the output of the attention module at layer $l$ be represented by $a^{l}$, and the output of the FFN module be represented by $m^{l}$. As the vectors computed in the attention and MLP modules get added back to the residual stream at each layer, the residual stream represents a summation of an increasing number of vectors as we go deeper into the model. A non-recursive formula for the output of the transformer just before unembedding and final layernorm is shown below:
@@ 6.1 Explaining Loss of Downstream Performance due to Norm-Growth
Here, $L$ represents the total number of layers in a model and $h^{L}$ represents the residual stream after the final layer. Thus, the output vector at the final layer is a summation of the outputs of individual self-attention and MLP sub-modules.
@@
Now, if the norm of the edited MLP matrix grows as disproportionately as shown in Figure 1, the norm of the vectors that are produced from those edited MLP sub-modules will also grow. This means that the norms of the vectors $m^{l}$ in the summation corresponding to the edited layers will grow substantially. As the norm of a few vectors in the summation grows, these vectors will begin to dominate the sum. Proof for this is shown in Appendix D.1, where we show that if the norm of a vector in a summation grows, the overall sum effectively tends towards that vector.
@@
We also show this effect empirically. The growing norm of activation vectors produced by edited layers after editing Llama3-8B can be seen in Figure 4. After editing using MEMIT, which edits layers 4-8 for 10,000 sequential edits (edited layers are shown in red color on the x-axis), we see a drastic increase in the norm of activation vectors produced by edited layers. For example, the activation vectors produced by layer 8 account for almost 40% of the total norm, and vectors produced by all edited layers account for 85% of the total. To emphasize how extreme this is, the residual stream for Llama-3-8B is made up of a summation of 65 vectorsLlama-3-8B has 32 layers, each contributing two vectros $a^{l}$ and $m^{l}$ to the residual stream along with the input vector., and 4 out of the 65 vectors coming from the edited layers account for 85% of the total norm. The norm-contributions for an unedited model can be seen in Figure 9(a) (appendix), where the four edited layers account for less than 4% of the overall norm. Note that this effect cannot be mitigated by LayerNorm or RMSNorm, which just normalize the incoming vectors Brody et al. (2023), whereas the norm growth of edited layers changes the content of the final hidden representation by making it consist of mostly the output of the edited layers.
@@
This gives us a crucial insight into why the growing norm of edited matrices leads to a loss of downstream performance. As the norms of the edited matrices increase, and as a consequence, the norm of the activation vectors produced from those matrices also increases, the activation vector at the final layer ($h^{L}$) is dominated by the output of edited layers, thus giving the edited layers a larger authority over the final representations. This allows edited layers to override the information produced from other parts of the model, possibly also helping make successful edits. However, as the norm of the edited matrix continues to grow when we edit it sequentially, suddenly the final representations begin to largely be composed of the outputs of only the edited layers. This makes it impossible for the model to account for the information processed from other sub-modules, which might be important to perform other downstream or unrelated tasks.
6.2 Introducing Norm Constraint
@@
In the above discussion, we show how growing norm of edited matrices is detrimental to the functioning of edited models and hypothesize how it causes loss of downstream performance. To test these conclusions, we propose to add an additional term to the editing objective to control this norm growth. Thus, we augment the MEMIT objective with a Frobenius norm-constraint:
@@
Figure 4: The figure shows the proportion of contribution of activation vectors from each sub-module to the residual stream. Edited layers are shown in red.
The original weight matrix is represented by $W_{0}$ and $\hat{W}$ represents the edited matrix. The above objective has a closed-form solution as shown below (proof, Appendix F):
@@
Table 3 and Figure 5 show the editing and downstream performance after 10,000 sequential edits using the norm-constraint. Compared to baseline methods such as MEMIT, AlphaEdit, as well as regularization approaches such as PRUNE and RECT, Frobenius norm-constraint version of MEMIT (represented by MEMIT + NC) produces consistent and substantial improvements in editing and downstream performance metrics. Specifically, we see in Table 3 that MEMIT + NC outperforms all previous models on all editing metrics except for the generation entropy metric. This shows that explicitly controlling norm growth is more effective than prior regularization methods like PRUNE and RECT, particularly when editing at scale. When looking at downstream performance in Figure 5, MEMIT + NC also outperforms all prior editing methods when looking for Llama-2-7B, and all methods except AlphaEdit for Llama3-8B. Moreover, the observed gains in downstream performance support our hypothesis that uncontrolled norm growth in edited layers leads to downstream degradation.
Method Edit Score Paraphrase Score Neighborhood Score Overall Score Generation Entropy Llama2-7B Llama3-8B Llama2-7B Llama3-8B Llama2-7B Llama3-8B Llama2-7B Llama3-8B Llama2-7B Llama3-8B MEMIT 81.04 49.68 64.67 49.29 60.95 51.31 67.86 50.08 442.59 373.48 PRUNE 70.80 49.38 62.11 49.63 51.86 51.09 60.60 50.02 280.83 340.22 RECT 82.42 63.17 66.84 56.92 67.39 52.89 71.54 57.36 549.35 588.39 AlphaEdit 61.10 72.67 55.86 63.44 53.75 52.90 56.74 61.95 540.92 465.81 AlphaEdit w/o NC 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 434.17 560.92 MEMIT + LTI + NC 54.41 50.76 51.87 48.89 52.05 51.17 52.75 50.25 516.24 575.59 MEMIT + NC 87.89 85.72 79.10 77.08 59.72 58.45 73.59 71.86 517.86 367.46 MEMIT + MPES + NC 91.09 87.54 78.02 77.97 59.78 59.30 74.03 72.97 545.72 536.43
Table 3: Editing performance of our approach when compared to baseline MEMIT, AlphaEdit and MEMIT regularization method such as PRUNE and RECT.
@@
Figure 5: Average downstream performance for during sequential editing with compared to baseline of MEMIT and addition of MPES and Norm-Constraint (NC).
Additionally, we want to note that AlphaEdit inherently has an explicit norm-constraint in its optimization objective. To present the importance of norm-constraint during large scale editing, we remove this term from the objective and evaluate the method, as shown in Table 3. To our surprise, we find that AlphaEdit completely fails without the norm constraint, as shown in the row ‘AlphaEdit w/o NC’ in Table 3 . Specifically, after 10,000 edits, the model effectively collapes, where Llama2-7B outputs <unk> and Llama3-8B loops on "!" and all editing scores drop to zero. We also illustrate the unbound growth of norm in Figures 26(a) and 26(b) in Appendix L. These results support our analysis in section 6.1 and re-emphasize that managing norm-growth is essential for stable and successful large-scale editing.
6.3 Combining MPES and Norm Constraint
In this section, we combine MPES, which intervenes in the gradient-descent step, with explicit Frobenius norm constraint, which regularizes the weight-update step in MEMIT. This combined approach, MEMIT + MPES + NC, achieves the strongest results across both editing score and downstream metrics. As shown in Table 3, it consistently improves the consistency of editing scores between models, with the largest gains observed on Llama3-8B. More importantly, Figures 5(a) and 5(b) show that the combination of the two regularization methods beautifully preserves downstream task performance with great effectiveness when compared to other algorithms, even after 10,000 sequential edits.
These findings suggest that MPES and norm-constraint address complementary failure modes during large-scale sequential editing. Their combination provides regularization methods that enable scalable knowledge editing while having minimal loss of downstream task performance. The generalizability of these findings can be seen by the compatibility of MPES with other knowledge editing methods, and the neccesity and importance of managing norm-growth in both MEMIT and AlphaEdit.